OpenStudy (anonymous):

how do i find (f/g)(x) for each f(x) and g(x) f(x)=x / (x+1) g(x)= (x^2)-1

OpenStudy (anonymous):

Well basically you take all of f(x) and put it over g(x) and simplify

OpenStudy (anonymous):

\[\Large \frac{\frac{x}{x+1}}{x^2-1}\]

OpenStudy (anonymous):

You know how to divide two fractions and simplify, correct?

OpenStudy (anonymous):

to find (f/g)(x)..put the value of g(x) at the place of every x in f(x) and similarly to find g(f(x) put the value of f(x) at the place of every x in g(x)

OpenStudy (anonymous):

Actually this is asking for (f/g)(x) not f(g(x)) or g(f(x)), this is similar to (f+g)(x) where you would just add the two functions together, but not replace every x value with one function and place it in the other f(g(x)) or g(f(x)) is not the same method as (f+g)(x) (f-g)(x) (f/g)(x) and (fg)(x)

OpenStudy (anonymous):

you got this?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

i thought that might be the case lets go slow

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

\[f(x)=\frac{x}{x+1}\\ g(x)=x^2-1\]right?

OpenStudy (anonymous):

that makes \[\frac{f(x)}{g(x)}=\frac{\frac{x}{x+1}}{x^2-1}\] as a first step, i..e put \(f\) on top of \(g\) ok with that?

OpenStudy (anonymous):

then do write it as one fraction, just like with numbers, put the denominators in the denominator and get \[\frac{f(x)}{g(x)}=\frac{x}{(x+1)(x^2-1)}\]

OpenStudy (anonymous):

and that is finished

OpenStudy (anonymous):

oh wow, that didnt look hard as i thought it would be. thank you for your help sir!

OpenStudy (anonymous):

yw (sir, i like that...)

OpenStudy (anonymous):

:D